OpenStudy (anonymous):

the positions of three particles are as follows x=3t+2.5t^2 x=5/2t(2+t) x=6 + 3t-2.5t^2 relative motion is uniform between (i) and (ii) (ii) and (iii) (i) and (iii) all of these

OpenStudy (inkyvoyd):

i and iii?

OpenStudy (anonymous):

inkyvoyd why did u block me..

OpenStudy (anonymous):

the answer is a

OpenStudy (inkyvoyd):

I didn't block you 0.o

OpenStudy (anonymous):

ohhh sorry>>

OpenStudy (inkyvoyd):

x=5/2t(2+t) is that t under the 5?

OpenStudy (anonymous):

no it is 5t/2

OpenStudy (anonymous):

did u get the answer inkyvoyd

OpenStudy (inkyvoyd):

so, t(5+(5t)/2)=(5t^2)/2+5t

OpenStudy (inkyvoyd):

I'd say that it's i and ii because they have the same coefficient for the term of second degree (ax^2)

OpenStudy (anonymous):

what u r saying the question is to find the relative motion is uniform in

OpenStudy (inkyvoyd):

Well, because the term of the second degree is the same, their velocities are both the same.

OpenStudy (anonymous):

the third one also has degree 2

OpenStudy (anonymous):

then......

OpenStudy (inkyvoyd):

Do you know how to tak the derivative?

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

Right now the equations we have describe displacement. The derivative of displacement is velocity. The derivative of that is acceleration. relative motion is uniform if they are the same.(the accel)

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

x=3t+2.5t^2 a=5 x=5/2t(2+t) a=5 x=6 + 3t-2.5t^2 a=-5

OpenStudy (inkyvoyd):

Since i and ii have the same accleration, that is the answer.

OpenStudy (anonymous):

thanxxx

OpenStudy (anonymous):

next question !!!

OpenStudy (inkyvoyd):

Alternative way is to look at it this way d=vt+(1/2)at^2 compare the a, and if it is the same, then you know they are the same.